OpenStudy (anonymous):

Find the sum. three-fourths plus one-half one-fourth two-thirds one and one-fourth 1

OpenStudy (anonymous):

one and one-fourth \[\frac{ 1 }{ 2 }\] \[\frac{ 3 }{ 4 }\] in order to do this you need common denominators so youd multiply the one half by 2 and thatll be \[\frac{ 2 }{ 4 }\] add it with three forths and you get five forths reduce it to get one and one forth so the answer is C